OpenStudy (anonymous):

Alinn flipped a coin 20 times. The coin landed heads up 12 times and tails up 8 times. Part A: Based on the results, what is the experimental probability of the coin landing heads up? Show your work. (5 points) Part B: What is the theoretical probability of the coin landing heads up? Show your work. (5 points)

OpenStudy (anonymous):

@horsegirl27 @hotguy @iGreen @michaeljordan23 @kingalex123 @satellite73

OpenStudy (anonymous):

second part should be easy right? coin has two sides and is fair

OpenStudy (anonymous):

oh yeah part b im okay with c:

OpenStudy (anonymous):

what do you think that is as a probability ?

OpenStudy (anonymous):

for part b?

OpenStudy (anonymous):

first part is "relative frequency": favorable over total

OpenStudy (anonymous):

i only see A and B B is clearly \(\frac{1}{2}\)

OpenStudy (anonymous):

so for part a i put relative frequency

OpenStudy (anonymous):

12/20 is the "relative frequency"

OpenStudy (anonymous):

thats what i would put c:

OpenStudy (anonymous):

????????

OpenStudy (anonymous):

i would probably reduce the fraction

OpenStudy (anonymous):

\[\frac{12}{20}=\frac{3}{5}\]

OpenStudy (anonymous):

good idea